OpenStudy (anonymous):

9=27k/25 solve for k

OpenStudy (anonymous):

how do you solve for k?

OpenStudy (anonymous):

k= 9*25/27

OpenStudy (anonymous):

\[9=\frac{27k}{25}\]\[9(25)=27k\]\[\frac{9(25)}{27}=k\]